# MSPipe: Efficient Temporal GNN Training via Staleness-aware Pipeline

### v1

@@ Abstract.
copyright: none
@@ 1. Introduction
Many real-world graphs exhibit dynamic characteristics, with nodes and edges continuously evolving over time, such as temporal social networks (Rozenshtein and Gionis, 2019; Wei and Carley, 2015) and temporal user-item graphs in recommendation systems (Ma et al., 2020; Ye et al., 2020). Previous attempts to model such dynamic systems have relied on static graph representations, which overlook their temporal nature (Zhang et al., 2019, 2020; Nguyen et al., 2018). Recently, temporal graph neural networks (TGNNs) have been developed to address this limitation. TGNNs are designed to incorporate time-aware information, learning both structural and temporal dependencies. Consequently, TGNNs facilitate more accurate and comprehensive modeling of dynamic graphs (Rossi et al., 2021; Wang et al., 2021; Kumar et al., 2019; Trivedi et al., 2019; Xu et al., 2020; Zhang et al., 2023; Cong et al., 2023; Sankar et al., 2020).
@@ 3. MSPipe framework
Table 1. Training time breakdown of TGN model. Profiling setup and other models’ statistics are in Appendix C.
@@ 3.1. MSPipe mechanism
Significant memory operations overhead. We consider a 5-stage abstraction of memory-based TGNN training, i.e., graph sampling, feature fetching, memory fetching, MTGNN training, and memory update. We conduct detailed profiling of the execution time of each stage, with time breakdown shown in Table 1,  5, and  6. Memory operations incur substantial overhead ranging from 36.1% to 58.6% of one iteration training time for different MTGNN models, while sampling and feature fetching do not, due to the 1-layer MTGNN structure. In Figure 1(b), memory fetching depends on memory vectors updated at the end of the last iteration, and has to wait for the relatively long MTGNN training and memory updating to finish
@@
$m^{(i)}_{v}$ $=\mathop{msg}\Big{(}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\tilde{s}^{(i-k)}_{v}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\tilde{s}^{(i-k)}_{u}},y_{uv}(t),\Delta t\Big{)}$
(2) ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\tilde{s}^{(i)}_{v}}$ $=\mathop{mem}\Big{(}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\tilde{s}^{(i-k)}_{v}},m^{(i)}_{v}\Big{)}$
$h^{(i)}_{v}$ $=\mathop{emb}\Big{(}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\tilde{s}^{(i)}_{v},\ \tilde{s}^{(i)}_{u}}\ |\ u\in\mathcal{N}(v)\Big{)}$
@@ 3.2. Stall-free Minimal-staleness Pipeline
Figure 8. Distribution of $\Delta t$ in WIKI dataset. Other datasets can be found in Appendix E.2.5
@@ 5. Experiments
We conduct experiments to evaluate the proposed framework MSPipe, targeting at answering the following research questions:
@@ 5.1. Experiment settings
Training settings. To ensure a fair comparison, we used the same default hyperparameters as TGL, including a learning rate of 0.0001, a local batch size of 600 (4000 for the GDELT dataset), and hidden dimensions and memory dimensions of 100. We train each dataset for 100 epochs, except for GDELT, which was trained in 10 epochs. We sampled the 10 most recent 1-hop neighbors for all datasets and constructed mini-batches with an equal number of positive and negative node pairs for sampling and subgraph construction during training and evaluation. The experiments are conducted under the transductive learning setting and we use average precision for evaluation metrics.
@@ 5.2. Expedited Training While Maintaining Accuracy
The results in Table 2 show that MSPipe improves the training throughput while maintaining high model accuracy. AP in the table stands for average model precision evaluated on the test set. For a more comprehensive analysis of various batch sizes, we provide detailed experiments in Appendix E.4.
@@
Scalability. Figure 9 presents the training throughput with different numbers of GPUs and the full results can be found in Appendix E.2.2. MSPipe achieves not only consistent speed-up but also up to 83.6% scaling efficiency on a machine, which is computed as the ratio of the speed-up achieved by using 4 GPUs to the ideal speed-up, outperforming other baselines. We also scale TGN training on GDELT to two machines with eight GPUs in Figure 9(b). Without explicit optimization for inter-machine communication, MSPipe still outperforms the baselines and exhibits better scalability.
((a)) LASTFM
((b)) GDELT
Figure 9. Scalability of training TGN.
@@ 5.4. Stall-free Minimal Staleness Bound
To further validate that MSPipe can find the minimal staleness bound without delaying the MTGNN training stage, we conduct a comparative analysis of accuracy and throughput between the minimal staleness bound computed by MSPipe and other different staleness bounds $k$. The results, depicted in Figure 11 and Figure 20 in Appendix E.2.4 (due to space constraints), consistently demonstrate that MSPipe achieves the highest throughput while maintaining the best accuracy compared to other staleness bound options. Additionally, the computed minimal staleness bounds for various datasets range from 2 to 4, providing further evidence for the necessity of accurately determining the minimal staleness bound rather than relying on random selection. Note that $k=1$ represents the baseline method of TGL without applying staleness.
((a)) WIKI
((b)) LASTFM
((a)) MOOC
((b)) GDELT
Figure 11. Throughput and AP on different staleness bound
@@ 5.5. Staleness Mitigation Mechanism
Hyperparameter analysis. We examine the effect of hyperparameter $\lambda$ on test accuracy, as depicted in Figure 14. We find that mitigating staleness with a larger $\lambda$ ($>0.8$) results in better model performance than TGL’s results, indicating that we should retain more of the original stale memory representations and apply a small portion of mitigation from their similar ones.
Figure 12. Staleness error comparison on TGN
@@
5.6. GPU memory and utilization
We present an analysis of the memory overhead associated with MSPipe, as staleness-based strategies generally require additional memory to enhance training throughput. Unlike other asynchronous training frameworks (Chen et al., 2022; Li et al., 2018; Wan et al., 2022; Peng et al., 2022) that introduce staleness during DNN or GNN parameter learning, MSPipe only introduces staleness within the memory module to break temporal dependencies. Each subgraph is executed sequentially, resulting in no additional hidden states during MTGNN computation. The extra memory consumption in MSPipe comes from the prefetched subgraph, including node/edge features and memory vectors. We provide a detailed analysis to determine the upper bound of this additional memory overhead, assuming maximum neighbor size for all nodes (i.e., $\mathcal{N}=10$) (see Appendix E.6 for a comprehensive analysis). Additionally, we measure the actual memory consumption using the torch.cuda.memory_summary() API in our experiments across all models and datasets. Table 4 shows that the observed additional memory usage in MSPipe aligns with our analyzed upper bound. Moreover, we compare the additional memory cost with the GPU memory size in Table 4, demonstrating that it constitutes a relatively small proportion (up to 2.92%) of the modern GPU’s capacity.
Table 4. Additional memory overhead of TGN when applying staleness. Other models can be found in Appendix E.6
@@ 6. Related Works
Asynchronous Distributed Training. Many studies advocate asynchronous training with staleness for DNN and static GNN models. For distributed DNN training, previous works  (Li et al., 2018; Chen et al., 2022; Recht et al., 2011; Ho et al., 2013; Dai et al., 2018) adopt stale weight gradients on large model parameters to eliminate communication overhead, while GNN models typically have much smaller sizes. For static GNN training, PipeGCN (Wan et al., 2022) and Sancus (Peng et al., 2022) introduce staleness in node embeddings under the full-graph training paradigm. Although these methods are effective for static GNNs, their effect is limited when applied to MTGNNs, from three aspects: 1) they focus on full graph training and apply staleness between multiple GNN layers to overlap the significant communication overhead with computation. In MTGNN training, the communication overhead is relatively small due to subgraph sampling and the presence of only one GNN layer. 2) all previous GNN training frameworks simply introduce a pre-defined staleness bound without explicitly analyzing the relationship between model quality and training throughput, potentially leading to sub-optimal parallelization solutions; 3) the unique challenges arising from the temporal dependency caused by memory fetching and updating in MTGNN training have not been adequately addressed by these frameworks. Therefore, these asynchronous training frameworks for DNN and static GNN are not suitable for accelerating MTGNNs. A detailed analysis of the related work can be found in Appendix B.
@@ Appendix B MORE DISCUSSION ON THE RELATED WORK
  3. (3)
    Bottlenecks: In full-graph training scenarios, such as PipeGCN (Wan et al., 2022) and Sancus (Peng et al., 2022), the main bottleneck lies in communication between graph partitions on GPUs. Due to limited GPU memory, the graph is divided into multiple parts, leading to increased communication time during full graph training. Therefore, these methods aim to optimize the communication-computation overlap to improve training throughput. In contrast, in MTGNN training, the main bottleneck stems from maintaining the memory module on the CPU and the associated challenges of updating and synchronizing it with CPU storage across multiple GPUs. MSPipe focuses on addressing this specific bottleneck. Furthermore, unlike full graph training where the entire graph structure needs to be stored in the GPU, MTGNN adopts a sampling-based subgraph training approach. As a result, the communication overhead in MTGNN is significantly smaller compared to full graph training.
  4. (4)
@@ E.2.2. Scalability results on all datasets
((a)) REDDIT
((b)) WIKI
((c)) MOOC
((d)) LASTFM
((e)) GDELT
@@
((a)) REDDIT
((b)) WIKI
((c)) MOOC
((d)) LASTFM
@@
((a)) REDDIT
((b)) WIKI
((c)) MOOC
((d)) LASTFM
@@
((a)) REDDIT
((b)) WIKI
((c)) MOOC
((d)) LASTFM
@@
((a)) REDDIT
((b)) WIKI
((c)) LASTFM
((d)) GDELT
@@ E.2.4. Comparison between different staleness bound
((a)) REDDIT
((b)) MOOC
((c)) LASTFM
((d)) GDELT

### v2

@@ Abstract.
Temporal Graph Neural Networks; Distributed Training; Efficient Training; Minimal Staleness Bound
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671844isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Artificial intelligenceccs: Information systems Information systems applications
@@ 1. Introduction
Many real-world graphs exhibit dynamic characteristics, with nodes and edges continuously evolving over time, such as temporal social networks (Rozenshtein and Gionis, 2019; Wei and Carley, 2015) and temporal user-item graphs in recommendation systems (Ma et al., 2020; Ye et al., 2020). Previous attempts to model such dynamic systems have relied on static graph representations, which overlook their temporal nature (Zhang et al., 2019, 2020; Nguyen et al., 2018; Su and Marbach, 2022; Su et al., 2024a). Recently, temporal graph neural networks (TGNNs) have been developed to address this limitation. TGNNs are designed to incorporate time-aware information, learning both structural and temporal dependencies. Consequently, TGNNs facilitate more accurate and comprehensive modeling of dynamic graphs (Rossi et al., 2021; Wang et al., 2021; Kumar et al., 2019; Su and Wu, 2023; Trivedi et al., 2019; Xu et al., 2020; Zhang et al., 2023; Cong et al., 2023; Sankar et al., 2020; Su et al., 2024b, 2023).
@@ 3. MSPipe framework
Table 1. Training time breakdown of TGN model.
@@ 3.1. MSPipe mechanism
Significant memory operations overhead. We consider a 5-stage abstraction of memory-based TGNN training, i.e., graph sampling, feature fetching, memory fetching, MTGNN training, and memory update. We conduct detailed profiling of the execution time of each stage, with time breakdown shown in Table 1. Memory operations incur substantial overhead ranging from 36.1% to 58.6% of one iteration training time for different MTGNN models, while sampling and feature fetching do not, due to the 1-layer MTGNN structure. In Figure 1(b), memory fetching depends on memory vectors updated at the end of the last iteration, and has to wait for the relatively long MTGNN training and memory updating to finish
@@
$m^{(i)}_{v}$ $=\mathop{msg}\Big{(}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{s}^{(i-k)}_{v}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{s}^{(i-k)}_{u}},y_{uv}(t),\Delta t\Big{)}$
(2) ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{s}^{(i)}_{v}}$ $=\mathop{mem}\Big{(}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{s}^{(i-k)}_{v}},m^{(i)}_{v}\Big{)}$
$h^{(i)}_{v}$ $=\mathop{emb}\Big{(}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{s}^{(i)}_{v},\ \tilde{s}^{(i)}_{u}}\ |\ u\in\mathcal{N}(v)\Big{)}$
@@ 3.2. Stall-free Minimal-staleness Pipeline
Figure 8. Distribution of $\Delta t$ in WIKI dataset. Other datasets follow a similar power-law distribution.
@@ 5. Experiments
We conduct experiments to evaluate the proposed framework MSPipe, targeting answering the following research questions:
@@ 5.1. Experiment settings
Training settings. To ensure a fair comparison, we used the same default hyperparameters as TGL, including a learning rate of 0.0001, a local batch size of 600 (4000 for the GDELT dataset), and hidden dimensions and memory dimensions of 100. We train each dataset for 100 epochs, except for GDELT, which was trained in 10 epochs. We sampled the 10 most recent 1-hop neighbors for all datasets and constructed mini-batches with an equal number of positive and negative node pairs for sampling and subgraph construction during training and evaluation. The experiments are conducted under the transductive learning setting and we use average precision for evaluation metrics. For a more comprehensive analysis of various batch sizes, we provide detailed experiments in Appendix E.4.
@@ 5.2. Expedited Training While Maintaining Accuracy
The results in Table 2 show that MSPipe improves the training throughput while maintaining high model accuracy. AP in the table stands for average model precision evaluated on the test set.
(a) LASTFM
(b) GDELT
Figure 9. Scalability of training TGN.
@@
Scalability. Figure 9 presents the training throughput with different numbers of GPUs on LastFM and GDELT datasets. MSPipe achieves not only consistent speed-up but also up to 83.6% scaling efficiency on a machine, which is computed as the ratio of the speed-up achieved by using 4 GPUs to the ideal speed-up, outperforming other baselines. We also scale TGN training on GDELT to two machines with eight GPUs in Figure 9(b). Without explicit optimization for inter-machine communication, MSPipe still outperforms the baselines and exhibits better scalability.
@@ 5.3. Preserving Convergence Rate
(a) WIKI
(b) LASTFM
(a) MOOC
(b) GDELT
Figure 11. Throughput and AP on different staleness bound
@@ 5.4. Stall-free Minimal Staleness Bound
To further validate that MSPipe can find the minimal staleness bound without delaying the MTGNN training stage, we conduct a comparative analysis of accuracy and throughput between the minimal staleness bound computed by MSPipe and other different staleness bounds $k$. The results, depicted in Figure 11, consistently demonstrate that MSPipe achieves the highest throughput while maintaining the best accuracy compared to other staleness bound options. Additionally, the computed minimal staleness bounds for various datasets range from 2 to 4, providing further evidence for the necessity of accurately determining the minimal staleness bound rather than relying on random selection. Note that $k=1$ represents the baseline method of TGL without applying staleness.
@@ 5.5. Staleness Mitigation Mechanism
Table 4. Additional memory overhead of TGN when applying staleness. Other models can be found in Appendix E.6.
@@
Hyperparameter analysis. We examine the effect of hyperparameter $\lambda$ on test accuracy, as depicted in Figure 14. The dashed horizontal lines in the figure denote the AP from the TGL baseline for comparison. We find that mitigating staleness with a larger $\lambda$ ($>0.8$) results in better model performance than TGL’s results, indicating that we should retain more of the original stale memory representations and apply a small portion of mitigation from their similar ones. Notably, setting $\lambda$ to 1 causes MSPipe-S to revert to the standard MSPipe configuration, thereby omitting the staleness mitigation strategy entirely.
5.6. GPU memory and utilization
We present an analysis of the memory overhead associated with MSPipe, as staleness-based strategies generally require additional memory to enhance training throughput. Unlike other asynchronous training frameworks (Chen et al., 2022; Li et al., 2018; Wan et al., 2022; Peng et al., 2022) that introduce staleness during DNN or GNN parameter learning, MSPipe only introduces staleness within the memory module to break temporal dependencies. Each subgraph is executed sequentially, resulting in no additional hidden states during MTGNN computation. The extra memory consumption in MSPipe comes from the prefetched subgraph, including node/edge features and memory vectors. We provide a detailed analysis to determine the upper bound of this additional memory overhead, assuming maximum neighbor size for all nodes (i.e., $\mathcal{N}=10$). Additionally, we measure the actual memory consumption using torch.cuda.memory_summary() API in experiments across all datasets. Table 4 shows that the observed additional memory usage in MSPipe aligns with our analyzed upper bound. Moreover, we compare the additional memory cost with the GPU memory size in Table 4, demonstrating that it constitutes a relatively small proportion (up to 2.92%) of the modern GPU’s capacity.
@@ 6. Related Works
Asynchronous Distributed Training. Many studies advocate asynchronous training with staleness for DNN and static GNN models. For distributed DNN training, previous works  (Li et al., 2018; Chen et al., 2022; Recht et al., 2011; Ho et al., 2013; Dai et al., 2018; Barkai et al., 2019) adopt stale weight gradients on large model parameters to eliminate communication overhead, while GNN models typically have much smaller sizes. For static GNN training, PipeGCN (Wan et al., 2022) and Sancus (Peng et al., 2022) introduce staleness in node embeddings under the full-graph training paradigm. Although these methods are effective for static GNNs, their effect is limited when applied to MTGNNs, from three aspects: 1) they focus on full graph training and apply staleness between multiple GNN layers to overlap the significant communication overhead with computation. In MTGNN training, the communication overhead is relatively small due to subgraph sampling and the presence of only one GNN layer. 2) all previous GNN training frameworks simply introduce a pre-defined staleness bound without explicitly analyzing the relationship between model quality and training throughput, potentially leading to sub-optimal parallelization solutions; 3) the unique challenges arising from the temporal dependency caused by memory fetching and updating in MTGNN training have not been adequately addressed by these frameworks. Therefore, these asynchronous training frameworks for DNN and static GNN are not suitable for accelerating MTGNNs. A detailed analysis of the related work can be found in Appendix B.
@@ 7. Conclusion
Acknowledgements.
We would like to thank the anonymous reviewers and area chairs for their helpful comments. This work is supported in part by grants from Hong Kong RGC under the contracts HKU 17207621, 17203522 and C7004-22G (CRF)
@@ Appendix B MORE DISCUSSION ON THE RELATED WORK
  • Bottlenecks: In full-graph training scenarios, such as PipeGCN (Wan et al., 2022) and Sancus (Peng et al., 2022), the main bottleneck lies in communication between graph partitions on GPUs. Due to limited GPU memory, the graph is divided into multiple parts, leading to increased communication time during full graph training. Therefore, these methods aim to optimize the communication-computation overlap to improve training throughput. In contrast, in MTGNN training, the main bottleneck stems from maintaining the memory module on the CPU and the associated challenges of updating and synchronizing it with CPU storage across multiple GPUs (Zhong et al., 2023). MSPipe focuses on addressing this specific bottleneck. Furthermore, unlike full graph training where the entire graph structure needs to be stored in the GPU, MTGNN adopts a sampling-based subgraph training approach. As a result, the communication overhead in MTGNN is significantly smaller than full graph training.
@@ E.2.2. Scalability results on all datasets
(a) REDDIT
(b) WIKI
(c) MOOC
(d) LASTFM
(e) GDELT
@@
(a) REDDIT
(b) WIKI
(c) MOOC
(d) LASTFM
@@
(a) REDDIT
(b) WIKI
(c) MOOC
(d) LASTFM
@@
(a) REDDIT
(b) WIKI
(c) MOOC
(d) LASTFM
@@
(a) REDDIT
(b) WIKI
(c) LASTFM
(d) GDELT
@@ E.2.4. Comparison between different staleness bound
(a) REDDIT
(b) MOOC
(c) LASTFM
(d) GDELT